Car A and Car B leave origin at 2 pm. Car A travles south at 60 mph and Car B travles east at 80 mph. Att what time will they be 200 miles apart?
Answer by Paul(988)

Let the time be x
60x+80x=200
140x=200
x=200/140
x=10/7

Hence, afte 10/7 hours they will be 200 miles apart.
Paul.